OpenStudy (anonymous):

Paul needs to construct two corrals for his pigs and his cows. The animals will be allowed to drink from the river so no fence is needed along the river. Paul has 97 feet of fence to complete the project. What should the dimensions of each corral be in order maximize the total enclosed area in the two corrals. [ ________________l R l [ l Pigs l i l Y -[ --------------l v l [ l Cows l e l [ --------------l r l X a) Identify the restriction and the objective functions

OpenStudy (anonymous):

b) Use your functions from (a) above, to write the area of the corrals as a function of the width x A(x) =

OpenStudy (anonymous):

We know 3X + Y = 97

OpenStudy (anonymous):

Y = 97 - 3X substitute in for Area(x) = X*Y = X(97-3X) 97X-3X^2

OpenStudy (anonymous):

Is that right?